OpenStudy (anonymous):

how to solve dy/dx: cos*2X

OpenStudy (anonymous):

derivative of cos (chain 2x)

OpenStudy (anonymous):

sin2x*2

OpenStudy (anonymous):

Are you sure, Starica is getting ready to correct you

OpenStudy (anonymous):

awww i thought we r to find dy/dx :S

OpenStudy (anonymous):

you just derive the outside and inside ^_^ outside = cos inside 2x so dy/dx of cos = -sin and dy/dx of 2x = 2 soooo, dy/dx cos(2x) = -2sin(2x)

OpenStudy (anonymous):

o_o... but wasn't she right?

OpenStudy (anonymous):

no i missed the negative sign

OpenStudy (anonymous):

isn't it finding the derivative of cos(2x)?

OpenStudy (anonymous):

lol, it's alright :)

OpenStudy (anonymous):

i guessed it so

OpenStudy (anonymous):

you were close =P

OpenStudy (anonymous):

thanks:p

OpenStudy (anonymous):

np ^_^, do you get it now sandip?

OpenStudy (anonymous):

no its finding the derivative of cos^2x

OpenStudy (anonymous):

OH! so the question is: \[\frac{dy}{dx} \cos^2(x)\]?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

alright so we have 2 steps, first outside then inside, our outside here is the power and the inside is cos(x) so dy/dx = - 2cos(x) sin(x) ^_^

OpenStudy (anonymous):

but star ...dont u think it must b y= cos^2x n find dy/dx??

OpenStudy (anonymous):

yeah lol

OpenStudy (anonymous):

wasn't I right? o_o

OpenStudy (anonymous):

... Go by the definition of the chain rule. \[Let\ (cos\ x)^2 = f(g(x))\] What is our f(x), what is our g(x)?

OpenStudy (anonymous):

she/he doesn't get the chain rule >_< lol

OpenStudy (anonymous):

in a much simpler way do it this way: cos(x) . cos(x) u'v + uv' ^_^ and solve

OpenStudy (anonymous):

I hope this way helps though :)

OpenStudy (anonymous):

alright gtg, later ^_^

OpenStudy (anonymous):

he-she isn't going to be able to do any of the rest of their homework if they don't understand the chain rule. This is a good opportunity to learn the chain rule since you need it to solve this problem.

OpenStudy (anonymous):

What can I say. Star likes manual labor.

OpenStudy (anonymous):

lol, I'm just trying to get it clear for him/her ^_^?